# CompBench: Benchmarking Complex Instruction-guided Image Editing

### v2

@@ 5 Insights
In this section, we investigate the underlying factors that lead to varying performances among different models on our proposed CompBench, and offer perspectives on future research directions for the field of image editing.
Figure 5: Overall Model Performance.(a) Top 5 model performace in five major evaluation tasks. (b) Overall model performace across all tasks.
The Critical Role of MLLMs. Through systematic evaluation, we discover a strong correlation between architectural design and editing performance: multi-modal large language models (MLLMs) [18, 1, 20, 34, 13, 12, 40, 17] serve as a cornerstone for recent advances in this field. Specifically, for a fair comparison, we normalize and average the performance scores of all models across five major tasks, highlighting the top-5 models in Figure 5(a). Furthermore, we present the overall normalized results in Figure 5(b), aggregated across all tasks. The results reveal that Step1X-Edit [21] significantly outperforms others on complex instruction-editing tasks, closely followed by SmartEdit [14] and GoT [7]. Notably, while MGIE [9] ranks 10th out of 13 models overall, it demonstrates comparatively performance on implicit reasoning tasks (particularly background evaluation metrics) which require advanced reasoning capabilities. Interestingly, all these models, including Step1X-Edit, SmartEdit, GoT and MGIE, integrate MLLMs with diffusion models to better fuse textual and visual information. This integration enables these models to more accurately understand complex instructions and the overall visual context, ultimately driving superior editing performance in complex editing tasks.
The Importance of Reasoning Ability. In addition to incorporating MLLMs to enhance visual-text fusion, we have also recognized the significant impact of reasoning ability on editing performance. SmartEdit is trained on the reasoning segmentation dataset from LISA, which significantly enhances its reasoning capabilities and leads to outstanding results on multiple tasks. On the other hand, with the recent development of reasoning MLLMs, GoT introduces Chain-of-Thought (CoT) [36] into the editing process by leveraging MLLMs to generate reasoning chains. This approach further improves the model’s understanding of editing instructions and images, enabling more precise editing.
Our analysis reveals two critical insights that advance the field of complex instruction-guided image editing. Fisrt, MLLMs are indispensable for high-performance editing, as they uniquely bridge the gap between textual instructions and visual understanding. Second, multi-modal reasoning capabilities are foundational to achieving high-quality editing outcomes. These findings indicate that future research should focus on: (1) developing MLLM architectures specifically optimized for editing workflows rather than general vision-language tasks (2) advancing reasoning-aware training frameworks, such as reinforcement learning (RL)-based optimization of reasoning chains or utilizing reasoning data, to enhance editing precision and adaptability.
@@ 6 Conclusion
In this work, we introduce CompBench, the first large-scale benchmark specifically designed for comprehensive evaluation of instruction-guided image editing capabilities. Our meticulously constructed benchmark encompasses five major categories with nine specialized tasks targeting complex image manipulation scenarios, comprising over 3,000 high-quality image editing pairs with corresponding natural language instructions. We conduct extensive experimental evaluation across 13 state-of-the-art instruction-guided image editing models on all benchmark tasks to systematically assess the capabilities and limitations of contemporary editing systems and validate the efficacy of our evaluation framework. The experimental findings from CompBench not only reveal significant performance gaps in current models but also yield valuable insights that elucidate promising research directions for advancing next-generation image editing systems with enhanced reasoning abilities and fine-grained control.
Appendix A Implementation Details
A.1 Source Data Collection and Preprocessing
Figure 6: Prompts of Editing Evaluation.
The primary data collection and preprocessing pipeline has been described in section 3.2. For initial quality assessment of MOSE data, we employed four no-reference metrics (NIQE [42], MANIQA [39], MUSIQ [16], and CLIPIQA [33]). All images were systematically evaluated using these metrics, with the scores subsequently normalized and equally weighted to compute a composite value for filtering purposes. In terms of mask preprocessing, we decomposed multi-object masks into discrete single-object representations and utilized a multimodal large language model (specifically Qwen-VL [34]) to quantitatively assess mask continuity and occlusion levels, with the corresponding prompts illustrated in Figure 6. The comprehensive data preprocessing workflow underwent rigorous multi-round review and verification by a dedicated team of domain experts, thereby ensuring the exceptional quality and reliability of the final dataset.
A.2 Task-specific Data Generation Pipelines
Figure 7: Local Editing Pipeline.
Figure 8: Comparison of Different Blending Operations.
Figure 9: Action/Scene Spatial Editing Pipeline.
Figure 10: Complex Reasoning Pipeline.
Figure 11: Multi-editing Pipeline.
Due to the distinct characteristics of different tasks, we have designed specialized pipelines for data generation tailored to related task categories. The specific workflows for each pipeline are as follows:
Local Editing Pipeline for Object-Level Manipulations. As illustrated in Figure 7, given the preprocessed high-quality VOS dataset, we employ an inpainting model (PowerPaint [46]) to execute object removal based on precise object masks. The resultant outputs undergo rigorous evaluation and refinement through a Multimodal Large Language Model (MLLM) [18, 1, 20, 34] in conjunction with manual verification. For instruction generation, we provide the pre- and post-edited images alongside the corresponding masks, utilizing the MLLM to generate preliminary instructions, which are subsequently refined manually to ensure they accurately reflect the specific editing operations. To facilitate comprehensive evaluation, we systematically select and manually annotate captions for the mask regions before and after editing, thereby generating the definitive Object Removal dataset. For Object Addition data, we strategically reverse a subset of the Object Removal data and similarly employ MLLM complemented by manual annotations to construct the corresponding instructions. For Object Replacement, we implement a candidate target object generation approach, wherein the MLLM analyzes the before-and-after images and masks from Object Removal to propose plausible, diverse replacement objects that maintain contextual coherence with the scene’s characteristics. Following manual selection of appropriate objects, we execute object replacement using the inpainting model, with instructions formulated analogously to those in the aforementioned tasks.
To further enhance background consistency and overall quality of inpainting results, we implement a composite post-processing strategy that integrates dilation, Gaussian blurring, and Poisson blending. Specifically, we first dilate the edges of the target object in the post-editing image, employing a kernel size of 20. Subsequently, Gaussian blurring is applied to the dilated edges with a kernel size of 15 and a $\sigma_{X}$ value of 3. Finally, Poisson blending is applied between the original image and the modified object region in the post-editing image to achieve the definitive result. As illustrated in Figure 8, the unprocessed edited images (direct inpainting outputs) exhibit noticeable blurriness and diminished clarity. Initially, we explored a rudimentary blending approach, directly merging the edited region with the background from the original image. However, this methodology resulted in pronounced and conspicuous boundaries, compromising the overall blending quality. By incorporating dilation and Gaussian blurring, the boundary integration improved substantially, yet unnatural chromatic discrepancies between the edited region and the surrounding background persisted. To address this limitation, we further integrated Poisson blending, which generated significantly superior results, yielding natural and seamless integration without perceptible artifacts.
Action/Scene Spatial Editing Pipeline. For action editing and scene spatial editing tasks (including location and viewpoint editing), as illustrated in Figure 9, we strategically select pertinent data from the VOS dataset. Specifically, for action editing, we extract frames from the same video sequence where the background remains consistent, while the objects undergo motion transformations. For location editing, we identify frames from identical video sequences where the background maintains constancy, but the object exhibits positional displacement. For viewpoint editing, we select frames exhibiting perspectival variations within the video. The generation of editing instructions for these tasks follows a methodology analogous to the local editing pipeline, integrating MLLM capabilities with meticulous manual refinement.
Complex Reasoning Pipeline. As depicted in Figure 10, for complex reasoning data, we systematically select a subset of data from the original VOS dataset that demonstrates suitability for implicit contextual edit instructions, subsequently modifying these instructions through manual intervention to produce the definitive dataset.
Multi-Editing Pipeline. For multi-object/multi-turn editing illustrated in Figure 11, we composite results from the object removal subset within the VOS dataset, where multiple object masks coexist within the same video frame, to generate comprehensive multi-object removal outputs. Conversely, this approach establishes the foundation for multi-object addition. For multi-turn editing, the outcome from manipulating a single object constitutes an initial editing iteration, while the multi-object composite result represents a subsequent editing phase, culminating in the multi-turn removal outcome. Similarly, this procedural framework can be applied to multi-turn addition operations. All associated instructions undergo rigorous manual development and refinement to ensure precision and clarity.
A.3 Experiment Setting Details
Table 5: Configurations Details of Editing Models.
Method Configuration
InstructPix2Pix [2] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
MagicBrush [38] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
[…]
Smart-edit [14] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
MGIE [9] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
HQ-Edit [15] guidance_scale=7, num_inference_steps=30 image_guidance_scale=1.5
CosXL-Edit [30] guidance_scale=7.5, num_inference_steps=20 image_guidance_scale=1.5
UltraEdit [45] guidance_scale=7.5, num_inference_steps=50 image_guidance_scale=1.5
AnyEdit [41] guidance_scale=3, num_inference_steps=100 image_guidance_scale=3
SEED-X [10] guidance_scale=7.5, num_inference_steps=100 image_guidance_scale=1.5
GoT [7] guidance_scale=5.0, num_inference_steps=50 image_guidance_scale=1.0
Step1X-Edit [21] guidance_scale=6, num_inference_steps=28
Evaluation Configurations. We basically use the official settings of all models during evaluation. Specific configurations of each model are shown in Table 5.
GPU Usage. All experiments using the inpainting model to construct our benchmark were conducted on 8 NVIDIA A6000 GPUs (48 GB each). The same setup was used during evaluation to meet the memory and computation demands of all editing models.
Appendix B Capability Analysis of CompBench
Table 6: Task Competence Analysis. We identify six core competencies essential for complex image editing. The degree to which each competency is required by a given task is classified as Low (L), Medium (M), High (H), or Ultra High (UH, bolded for extraordinarily high requirements).
Task Visual Appearance Relation Complex Scene 3D Grounding Control Understanding Reasoning Consistency Geometry Object Removal H H H H UH L Object Addition H UH H H UH L Object Replacement H UH H H UH L Action Change M H H M UH H Location Change H M M H UH M Viewpoint Change M M M M UH UH Multi-turn Editing H UH H H UH M Multi-object Editing UH UH UH H UH H Implicit Reasoning H H UH UH UH H
To comprehensively characterize the requirements of complex image editing, we identify six core competencies that a sophisticated editing system must demonstrate: (1) Visual Grounding. the precise localization of target objects or regions; (2) Appearance Control. fine-grained manipulation of visual attributes such as color, texture, and illumination; (3) Relation Understanding. accurate modeling of semantic and spatial dependencies among objects; (4) Complex Reasoning. implicit logical deduction from contextual cues; (5) Scene Consistency. holistic preservation of spatial layout, occlusion patterns, and contextual coherence; and (6) 3D Geometry. understanding and manipulating three-dimensional structure and viewpoint.
The quantitative correspondence between individual tasks and the competencies they necessitate is summarized in Table 6. Local editing tasks place an exceptionally high premium on Scene Consistency, as seamless integration of the modified region is paramount. Multi-object Editing requires a balanced proficiency in Appearance Control, Relation Understanding, and Scene Consistency to effectively manage complex inter-object interactions. Viewpoint Editing uniquely depends on the 3D Geometry competency to facilitate perspective transformations, whereas Implicit Reasoning imposes stringent demands on Complex Reasoning to infer indirect editing intentions.
Overall, CompBench presents substantial challenges by requiring the simultaneous integration of these multi-dimensional competencies, thereby reflecting the intricacy of real-world image-editing scenarios.
Appendix C Cases of CompBench
Figure 12: Cases of Local Editing in CompBench.
Figure 13: Cases of Multi-editing in CompBench.
Figure 14: Cases of Action, Location and Viewpoint Editing in CompBench.
Figure 15: Cases of Local Editing Results.
Figure 16: More Cases of Local Editing Results.
Figure 17: Action, Location and Viewpoint Editing Examples.
Figure 18: Action, Location and Viewpoint Editing Examples.
Figure 19: Action, Location and Viewpoint Editing Examples.
In this section, we present additional exemplars from CompBench and comprehensive evaluation results. In Figure 12, we demonstrate representative instances of local editing operations (object addition, object removal, and object replacement) within our CompBench. In Figure 13, we illustrate selected cases of multi-turn and multi-object editing outcomes in our benchmark. In Figure 14, we showcase exemplary instances of action editing, location editing, and viewpoint editing capabilities within our benchmark.
We further present qualitative results from all evaluated models on our benchmarks. The comparative editing outputs across all models for local editing, multi-editing, and implicit reasoning tasks can be examined in Figures 15 and 16. The corresponding results for action editing, location editing, and viewpoint editing are displayed in Figures 17,18, and19. Detailed evaluation protocols and analytical discussions are presented in Section D.
Appendix D Evaluation Details
In this section, we delineate the comprehensive evaluation methodology and procedural framework employed in our assessment.
For tasks encompassing local editing, multi-editing, and implicit reasoning, we require models to modify foreground elements while maintaining background fidelity. Consequently, we evaluate editing performance from both foreground and background perspectives. Background consistency is quantitatively assessed utilizing PSNR, SSIM, and LPIPS metrics on background regions, while foreground evaluation incorporates both editing accuracy via CLIP image embedding similarity with ground truth exemplars and instruction adherence via CLIP Score with the local foreground caption. The foreground captions are meticulously annotated and validated, with representative examples illustrated in Figure 12.
To objectively quantify the quality of Action Editing, Location Editing, and Viewpoint Editing, which require substantial image manipulations, we devise an automated evaluation pipeline leveraging GPT-4o: for each source–edited image pair, the pipeline generates a reproducible and interpretable composite score ranging from 0 to 10, accompanied by a concise textual justification. Initially, we formulate three task-specific prompts as demonstrated in Figure 6, each emphasizing two to three criteria extracted from the fundamental task characteristics: action-execution correctness, preservation of non-target regions, and overall realism for Action; positional accuracy, occlusion consistency, and global harmony for Location; and plausibility of viewpoint transformation, geometric coherence, and detail preservation for Viewpoint. All prompts instruct GPT-4o to return a standardized JSON object of the form {"score": <0-10>, "reason": "<...>"}. Subsequently, each image pair is processed by GPT-4o precisely once with the decoding temperature fixed at 0, yielding deterministic and consequently fully reproducible results.
Appendix E Human Annotations
Annotation Details. To complete our benchmark development, we engaged four expert annotators and one quality assurance specialist to supervise all human verification processes throughout the benchmark construction. This comprehensive oversight encompassed filtering the initial datasets, validating generated images, and most critically, formulating precise instructions. During the instruction annotation phase, annotators were provided with the original images, edited images, and corresponding masks, and were tasked with crafting instructions based on this multimodal information. All annotators underwent extensive training, including exposure to exemplary cases, and received iterative feedback to progressively enhance their performance. To ensure annotation fidelity and inter-annotator consistency, the quality assurance specialist conducted systematic evaluations of all outputs. Significantly, all procedures were implemented in a double-blind framework to mitigate potential experimental biases.
Participant Disclosure and Consent. Annotators were comprehensively briefed regarding the purpose of the annotation task, their unconditional right to withdraw participation at any juncture, and the exclusive research-oriented utilization of their annotations. No personally identifiable information was collected during this process. All annotators provided documented informed consent prior to their participation in the study.
Compensation. Annotators received remuneration at a rate commensurate with or exceeding the local minimum wage standards.
Appendix F Statement of Limitation, Ethical Concern and Broader Impact
Limitation. A significant constraint of our methodology resides in the substantial computational overhead and procedural complexity of the integrated pipeline. The framework encompasses multiple sequential stages of model-based processing followed by comprehensive human evaluation, necessitating considerable computational resources and expert human intervention. This intricate design imposes reproducibility challenges, hampers scalability, and restricts the exhaustive coverage of the complete spectrum of compositional editing tasks. Furthermore, components leveraging large language models are inherently bounded by contemporary model limitations, potentially compromising performance on tasks demanding sophisticated reasoning or precise visual-linguistic integration.
Ethical Consideration. While our instruction-based image editing framework demonstrates robust capabilities, it elicits ethical considerations regarding potential misappropriation and privacy implications. The capacity to generate photorealistic manipulations may be exploited to create deceptive content, potentially exacerbating misinformation dissemination if inadequately regulated. Moreover, the utilization of real-world imagery introduces non-negligible risks of inadvertently exposing personally identifiable information, despite rigorous anonymization protocols. Ensuring responsible deployment necessitates implementing comprehensive safeguards, including content verification mechanisms, stringent data protection frameworks, and transparent operational guidelines.
Broader Impact. CompBench establishes a rigorous evaluation framework for instruction-guided image editing, facilitating systematic assessment of complex editing capabilities in multimodal large language models. This contribution will accelerate the development of reasoning-aware and controllable editing systems, enhance model performance across visual understanding and generation domains, and expand the practical applicability of large-scale models in diverse real-world contexts, including creative design processes, digital content production, and interactive artificial intelligence assistants.

### v4

@@ 5 Insights
In this section, we provide deeper insights and future directions based on model performances and failure analysis on our proposed CompBench.
The Critical Role of MLLMs.
We observe a strong correlation between architectural design and editing performance. As shown in Figure 6(a) and (b) (calculation details in supplementary), the top-performing models (Bagel [5], Qwen-Image-Edit [38], FLUX.1 Kontext [16], and Step1X-Edit [21]) mostly integrate multi-modal large language models (MLLMs). Models lacking MLLM integration frequently ignore instructions or edit wrong targets, indicating that standard CLIP-text alignment is insufficient for complex reasoning. Therefore, integrating MLLMs is indispensable for accurately interpreting intricate textual instructions and visual contexts.
Planner-Executor Misalignment.
Despite the advantages of MLLMs, our failure analysis reveals a widespread planner-executor misalignment. In high-density scenes, even when MLLMs (Planner) correctly identify targets, diffusion models (Executor) often fail at precise masking, causing background leakage. Unified architectures help mitigate this mismatch, as exemplified by Bagel’s joint learning of understanding and generation. Consequently, future research should focus on improving pixel-level grounding stability in cluttered scenes, moving beyond purely semantic understanding.
Reasoning Bottleneck.
Beyond basic semantic alignment, complex multi-modal reasoning is foundational for high-fidelity edits. Enhancing reasoning capabilities significantly boosts performance, evident in data-centric strategies (e.g., SmartEdit training on LISA [17]’s reasoning segmentation data) and method-centric designs (e.g., GoT introducing Chain-of-Thought [37] via MLLMs). Therefore, a crucial direction for future work involves continuously enhancing the reasoning capabilities of MLLMs through reasoning-aware training paradigms to ensure precise intent interpretation.
Geometric Hallucination.
Beyond semantics, we observe severe limitations in physical consistency. For complex spatial tasks like Action and Viewpoint editing, models frequently hallucinate distorted geometries. To address this issue, future frameworks need to incorporate 3D structural priors or geometric guidance into 2D-trained editors to maintain strict physical consistency during generation.
@@ 6 Conclusion
In this work, we introduce CompBench, the first large-scale benchmark designed for complex instruction-guided image editing. Our benchmark encompasses five major categories with nine specialized tasks comprising over 3,000 high-quality image editing pairs with corresponding instructions. We conduct extensive evaluation to systematically assess the capabilities and limitations of contemporary editing systems and validate our evaluation framework. Our findings not only reveal significant performance gaps in current models but also provide valuable insights to guide future research toward next-generation image editing systems.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (NO. 62572193), the Open Research Fund of the Key Laboratory of Advanced Theory and Application in Statistics and Data Science, Ministry of Education, and the Fundamental Research Funds for the Central Universities.